# Multi-Input Multi-Output Target-Speaker Voice Activity Detection For Unified, Flexible, and Robust Audio-Visual Speaker Diarization

### v2

@@ III-B Multi-Stage Training
To process multiple input and output types in a unified framework, we design a multi-stage training strategy to optimize the model progressively, as shown in Fig. 3. Meanwhile, we introduce two categories of modality masking techniques during the training, which enable the model to process audio-only, video-only, or audio-visual signals flexibly. The training process can be described as follows.
  • Stage 1: We copy and freeze the parameters of the pre-trained speaker embedding model to initialize the audio front-end extractor. Only audio and video-based output branches are adopted to train fully simulated data.
@@ III-B2 Data-Level Modality Masking
where $\vee$ denotes the element-wise logical OR operation. The $n$-th speaker’s ground truth labels should be valid as long as one of the $\mathbf{e}_{n}$ or $\mathbf{l}_{n}$is not zeroed. This way, the trained decoder can be compatible with uncertain speaker profiles.
@@ IV-A Datasets
The proposed MIMO-TSVAD framework is a multi-modal extension of our previous work [19], which supports audio-based, video-based, and audio-visual speaker diarization. We first verify its advantages on the VoxConverse [82] and DIHARD-III [83] datasets for audio-based speaker diarization. The VoxConverse dataset is an in-the-wild dataset with 20.29 hours of development set and 43.53 hours of test set collected from YouTube. We select the first 172 recordings (80%) of the original development set for training MIMO-TSVAD models. The last 44 recordings (20%) remain for validation. The DIHARD-III dataset is a multi-domain dataset with 34.15 hours of development set and 33.01 hours of evaluation set, including 11 complex scenarios (e.g., interview, clinical, restaurant). We select the first 203 recordings (80%) of the original development set for training MIMO-TSVAD models. The last 51 (20%) recordings remain for validation.
In addition, the Multi-Modal Information based Speech Processing (MISP) 2022 [37] dataset is adopted for audio-visual speaker diarization. The MISP 2022 dataset targets the Chinese home-TV scenario with 106.09 hours of training set, 2.6 hours of development set, and 3.12 hours of evaluation set. Audio-visual signals are synchronously captured by different devices. The single-channel microphone records the near-field ($<0.5m$) audio. The 2-channel microphone array and high-definition camera capture the middle-field ($1-1.5m$) data. The 6-channel microphone array and wide-angle camera capture the far-field ($3-5m$) data. Also, we utilize the NARA-WPE  [^2] https://github.com/fgnt/nara_wpe and SETK  [^3] https://github.com/funcwj/setk toolkits to do dereverberation and MVDR-based beamforming for the multi-channel audio signals, augmenting the diversity of training data. Under the MISP challenge rules, all fields can be used in model training and validation. Only far-field data is allowed for evaluation.
To benefit neural networks from large-scale training data, we create a simulated audio corpus by the VoxCeleb2 [84] dataset. As the speaker diarization task does not force the audio and video data to be semantically consistent, we cut the training set of the MISP 2022 [37] dataset into speech and non-speech segments as supplementary video corpus. Each simulated data is generated by mixing source utterances of 1-4 speakers from the audio-visual corpus. The average overlap ratio of simulated data is controlled to around 30%.
@@ IV-B1 Speaker Embedding Extraction
The ResNet-34 is utilized as the pattern extractor, where residual blocks have widths (number of channels) of $\left\{64,128,256,512\right\}$ with a total downsampling factor of 8. After adding the statistical pooling [42] layer, a linear projection layer outputs the 256-dim speaker embedding. The ArcFace ($s=32,m=0.2$[85] is used as the classifier. The model trained on the VoxCeleb2 [84] dataset achieves an equal error rate (EER) of 0.814% on the Vox-O [86] trial. Other training details are the same as [87]. The pre-trained model is used for speaker embedding extraction in the subsequent MIMO-TSVAD system. Also, the Speaker Alignment (SA) module employs the pre-trained speaker verification threshold of 0.3479 during inference.
@@ IV-B2 MIMO-TSVAD
The audio front-end ResNet-34 is initialized by the pre-trained speaker embedding model. We replace the original statistical pooling (SP) with segmental statistical pooling (SSP) [18] layer to achieve frame-level feature extraction. The video front-end ResNet18-3D has residual blocks with widths of $\left\{32,64,128,256\right\}$ to output 256-dim frame-level features, which are randomly initialized. The back-end encoder-decoder modules have 6 blocks sharing the same settings: 512-dim attentions with 8 heads and 1024-dim feed-forward layers with a dropout rate of 0.1. The kernel size of convolutions in Conformer blocks is 15, and the other implementation details are the same as [74].
@@
IV-C Training and Inference Details
All training data is split into fixed-length chunks by sliding window and normalized with a mean of 0 and a standard deviation of 1. The input acoustic features are 80-dim log Mel-filterbank energies with a frame length of 25 ms and a frameshift of 10 ms. The input lip videos are transformed into grayscale with a resolution of $88\times 8$ 88 and frames per second (FPS) of 25. Assume that the maximum number of speaker profiles (speaker capacity) is set to $C$. When speakers in a recording cannot reach $C$, empty speaker profiles are padded by zeros or speakers not appearing in the current chunk. This padding method aligns the dimension of batched training data and forces the model to distinguish valid and invalid speaker profiles. Meanwhile, all the input speaker profiles should be shuffled to make the model invariant to speaker order.
We implement the BCE loss and Adam optimizer to train the neural network. As described in Section III-B, the first training stage starts with a linear learning rate warm-up from 0 to 1e-4 in 2,000 iterations. From the second training stage, real data from the VoxConverse [82] and DIHARD-III [83] datasets is added to the simulated data at a ratio of 0.2. For experiments on the MISP 2022 [37] dataset, this ratio is adjusted to 0.5, which follows our previous challenge setting [36] to mitigate the large domain gap between simulated data (English) and test set (Chinese). In the last finetuning stage, the learning rate is decayed to 1e-5. Additive noise from Musan [92] and reverberation from RIRs [93] are applied for audio augmentation. For video augmentation, input lip videos undergo each item of the following procedures with a probability of 0.5: rotation with an angle range $[5,20]$; horizontal flipping; cropping with the scale range $[0.8,1]$; transformation of contrast, brightness, and saturation in the range $[-25,25]$. The training process takes around 200k iterations with a batch size of 32 on 8 $\times$ NVIDIA RTX-3090 GPUs.
@@ V-B Evaluation of Video-based and Audio-Visual Diarization
For the MISP 2022 evaluation set, reference VAD is provided according to the challenge’s track 1 rules. The LSTM-SC obtains the 29.30% DER, better than the AHC (30.02%) and VBx (33.37%). Thus, Table V illustrates the performance of different MIMO-TSVAD models initialized by the LSTM-SC result. To investigate the potential of audio-based, video-based, and mixed output methods, Systems S1-3 are individually trained for each inference stage. Experimental results show that the video-based System 2 obtains 15.01% DER, surpassing the audio-based System 1 with 23.35% DER. The improvement mostly comes from the speaker confusion (SF) error decreasing from 13.49% to 4.34%. Visual modality demonstrates a significant advantage in determining speaker identities as it does not suffer from the issue of overlapping speakers. Then, System 3 decreases the DER to 10.06% by introducing auxiliary audio information. About $3/4$ improvement (3.8% of 4.95% DER) is contributed by miss (MI) and false alarm (FA) rates. Compared with video signals of a limited 25 FPS, it can be seen that the audio modality can bring higher temporal precision for predicted utterance boundaries.
@@ V-C Evaluatioin of Robustness to Lip-Missing Scenarios
A speaker’s lip track may not always be available in real scenarios. To explore the impact of lip-missing problems on the MIMO-TSVAD framework, we newly simulate test data based on the original MISP 2022 evaluation set. Three lip-missing scenarios are created by removing the Lip-RoI randomly with different degrees, described as follows.
@@ VI Conclusions
This paper proposes a novel Multi-Input Multi-Output Target-Speaker Voice Activity Detection (MIMO-TSVAD) framework to tackle speaker diarization under complicated audio-visual data accessibilities. The model with jointly designed multi-stage training and inference strategies is compatible with different scenarios in a unified framework. Experiment results show that the MIMO-TSVAD framework performs well for audio-based, video-based, and audio-visual speaker diarization. It obtains new state-of-the-art DERs of 4.18% on the VoxConverse [82] test set, 10.10% on the DIHARD-III [83] evaluation set, and 8.15% on the MISP 2022 [37] evaluation set, respectively. Furthermore, the MIMO-TSVAD framework demonstrates strong robustness against lip-missing problems. In simulated scenarios with varying lip-missing degrees, it guarantees that the DERs of the audio-visual system are always no worse than the audio-only system. In the future, we will further improve the current approach regarding advanced lip extraction and clustering methods, better use of multi-channel audio, etc.

### v3

@@ III-B Multi-Stage Training
To process multiple input and output types in a unified framework, we design a multi-stage training strategy to optimize the model progressively, as shown in Fig. 3. Meanwhile, we introduce two categories of modality masking techniques during training, which enable the model to process audio-only, video-only, or audio-visual signals flexibly. The training process can be described as follows.
  • Stage 1: We copy and freeze the parameters of the pre-trained speaker embedding model to initialize the audio front-end extractor. Only audio and video-based output branches are adopted to train on fully simulated data.
@@ III-B2 Data-Level Modality Masking
where $\vee$ denotes the element-wise logical OR operation.
Since the masking probability is independent for all speakers, speaker embeddings or lip tracks of multiple speakers may be masked at a time. Because for the same speaker, we can only randomly mask either the speaker embedding or lip track simultaneously. If the $n$-th speaker is masked among all $N$ inputs, the model still generates an output of shape $N\times T^{\prime}$. As long as one of the $\mathbf{e}_{n}$ or $\mathbf{l}_{n}$ is available, the mixed output branch of the proposed model can work properly for the $n$-th speaker. This way, the trained decoder can be compatible with uncertain speaker profiles.
@@ IV-A Datasets
The proposed MIMO-TSVAD framework is a multi-modal extension of our previous work [19], which supports audio-based, video-based, and audio-visual speaker diarization. We first verify its advantages on the VoxConverse [81] and DIHARD-III [82] datasets for audio-based speaker diarization. The VoxConverse dataset is an in-the-wild dataset with 20.29 hours of development set and 43.53 hours of test set collected from YouTube. We select the first 172 recordings (80%) of the original development set for training MIMO-TSVAD models. The last 44 recordings (20%) remain for validation. The DIHARD-III dataset is a multi-domain dataset with 34.15 hours of development set and 33.01 hours of evaluation set, including 11 complex scenarios (e.g., interview, clinical, restaurant). We select the first 203 recordings (80%) of the original development set for training MIMO-TSVAD models. The last 51 recordings (20%) remain for validation.
[…]
We also create the simulated audio-visual dataset to benefit neural networks from large-scale training data. The VoxCeleb2 [83] dataset is adopted as the source audio corpus, which covers thousands of unique speaker identities. As the video contents of the VoxCeleb2 dataset are not available for downloading nowadays, we utilize the MISP 2022 [37] dataset as a supplementary video corpus. Although the audio and video signals are not semantically consistent regarding the dialogue content, the combination is still feasible because speaker diarization primarily focuses on recognizing speaker identities and voice activities rather than the speech content. During training, each simulated data is generated in an on-the-fly manner as follows.
  • Step 1: A speaker label is randomly selected from the audio corpus. The single-speaker utterance is created by alternately concatenating his or her source speech and silent (zero-padded) segments, where each segment length is sampled from a uniform distribution of 0-4 seconds.
  • Step 2: We cut the video corpus into active (speech) and inactive (silent) lip tracks. An active lip track is randomly selected and cropped into the same duration for each speech segment created in the first step. An inactive lip track is randomly selected and cropped into the same duration for each silent segment. Then, the extracted lip tracks are concatenated in the same order as the audio, creating the corresponding single-speaker lip track.
  • Step 3: For each simulation, audio-visual data of 1-4 speakers can be obtained by repeating the above steps. All single-speaker utterances are averaged as the mixed audio. No additional operations are required for the lip tracks, as they are already separate. The average overlap ratio of simulated data is estimated to be around 30%.
@@ IV-B1 Speaker Embedding Extraction
[…]
@@ IV-B2 MIMO-TSVAD
[…]
IV-C Training and Inference Details
All training data is split into fixed-length chunks by sliding window and normalized with a mean of 0 and a standard deviation of 1. The chunk length can be set to different values (e.g., 8 seconds, 16 seconds, 32 seconds). As the model only accepts a fixed-length chunk as input, the chunk lengths used for training and inference should always be consistent, but the chunk shift can vary flexibly. The subsequent experimental results investigate the impacts of different chunk lengths and shifts. Then, the input acoustic features are 80-dim log Mel-filterbank energies with a frame length of 25 ms and a frameshift of 10 ms. The input lip videos are transformed into grayscale with a resolution of $\times 88$ 88 and frames per second (FPS) of 25. Assume that the maximum number of speaker profiles (speaker capacity) is set to $C$. When speakers in a recording cannot reach $C$, empty speaker profiles are padded by zeros or speakers not appearing in the current chunk. This padding method aligns the dimension of batched training data and forces the model to distinguish valid and invalid speaker profiles. Meanwhile, all the input speaker profiles should be shuffled to make the model invariant to speaker order.
We implement the BCE loss and Adam optimizer to train the neural network. As described in Section III-B, the first training stage starts with a linear learning rate warm-up from 0 to 1e-4 in 2,000 iterations. From the second training stage, real data from the VoxConverse [81] and DIHARD-III [82] datasets is added to the simulated data at a ratio of 0.2. For experiments on the MISP 2022 [37] dataset, this ratio is adjusted to 0.5, which follows our previous challenge setting [36] to mitigate the large domain gap between simulated data (English) and test set (Chinese). In the last finetuning stage, the learning rate is decayed to 1e-5. Additive noise from Musan [87] and reverberation from RIRs [88] are applied for audio augmentation. For video augmentation, input lip videos undergo each item of the following procedures with a probability of 0.5: rotation with an angle range [5,20][5,20][ 5 , 20 ]; horizontal flipping; cropping with the scale range [0.8,1][0.8,1][ 0.8 , 1 ]; transformation of contrast, brightness, and saturation in the range [25,25][-25,25][ - 25 , 25 ]. The training process takes around 200k iterations with a batch size of 32 on 8 $\times$ NVIDIA RTX-3090 GPUs.
@@ V-A Evaluation of Audio-based Diarization
Notably, the DER results of MIMO-TSVAD models in Tables I and III differ slightly from those reported in the original Seq2Seq-TSVAD paper [19]. A clarification of the inference configurations is provided as follows.
  • First, for the VoxConverse dataset, this work does not apply any post-processing based on estimated VAD information, whereas [19] revises TS-VAD predictions using results from a separate diarization system. In this case, we observed that a finer-grained inference strategy, achieved by adjusting chunk lengths and chunk shifts, already leads to significant DER improvements. As a result, the best-performing MIMO-TSVAD model achieves competitive accuracy without requiring additional VAD-based refinement.
  • Second, for the DIHARD-III dataset, [19] directly reports the best DER under a specific configuration (chunk length = 16 s, VAD resolution = 10 ms, speaker capacity = 20). In contrast, this work systematically explores a wide range of inference setups to gain a deeper understanding of their impact on performance.
  • Third, the reported DERs in this work are averaged over three independent runs to mitigate randomness, whereas [19] reports single-run outcomes. These differences may account for the observed variations in DER.
@@ V-B Evaluation of Video-based and Audio-Visual Diarization
For the MISP 2022 evaluation set, reference VAD is provided according to the challenge’s track 1 rules. The LSTM-SC obtains the 29.30% DER, better than the AHC (30.02%) and VBx (33.37%). Thus, Table V illustrates the performance of different MIMO-TSVAD models initialized by the LSTM-SC result. To investigate the potential of audio-based, video-based, and mixed output methods, Systems S1-3 are individually trained for each inference stage. Experimental results show that the video-based System 2 obtains 15.01% DER, surpassing the audio-based System 1 with 23.35% DER. The improvement mostly comes from the speaker confusion (SC) error decreasing from 13.49% to 4.34%. Visual modality demonstrates a significant advantage in determining speaker identities as it does not suffer from the issue of overlapping speakers. Then, System 3 decreases the DER to 10.06% by introducing auxiliary audio information. About 3/43/43 / 4 improvement (3.8% of 4.95% DER) is contributed by miss (MI) and false alarm (FA) rates. Compared with video signals of a limited 25 FPS, it can be seen that the audio modality can bring higher temporal precision for predicted utterance boundaries.
@@ V-C Evaluatioin of Robustness to Lip-Missing Scenarios
A speaker’s lip track may not always be available in real scenarios. To explore the impact of lip-missing problems on the MIMO-TSVAD framework, we newly simulate test data based on the original MISP 2022 evaluation set. Three lip-missing scenarios are created using zeros to randomly mask the Lip-RoI for simulating off-screen data, described as follows.
@@
V-D Computing Efficiency
Table VII illustrates the computing efficiency of MIMO-TSVAD models trained in Table V. For each 8-second input data, several metrics for the different inference stages are presented as follows. First, the number of parameters for each stage is an essential indicator. Second, Floating-Point Operations (FLOPs) are used to measure the computational complexity. Then, we count the required GPU memory and the average time for inferring each input chunk, which is tested on the NVIDIA RTX-3090 GPU. Starting from the second stage, the use of visual modality significantly increases the computational load. The more multi-modal information the model uses, the greater the computational load is required. Improving the computing efficiency of multi-modal speaker diarization systems is still challenging.
TABLE VII: Computing Efficiency of MIMO-TSVAD models for each 8-second input data in different inference stages, regarding the number of parameters, Floating-Point Operations (FLOPs), GPU memory, and inference time.
[b] Stage Params (M) FLOPs (G) Memory (MB) Time (s) 1 76.56 151.80 494.49 0.0279 2 64.11 515.60 943.60 0.0499 3 85.78 667.01 1040.50 0.0617 4 153.72 669.16 1313.22 0.0797
@@ VI Conclusions
This paper proposes a novel MIMO-TSVAD framework to tackle speaker diarization under complicated audio-visual data accessibilities. The model with jointly designed multi-stage training and inference strategies is compatible with different scenarios in a unified framework. Experimental results show that the MIMO-TSVAD framework performs well for audio-based, video-based, and audio-visual speaker diarization. It obtains new state-of-the-art DERs of 4.18% on the VoxConverse [81] test set, 10.10% on the DIHARD-III [82] evaluation set, and 8.15% on the MISP 2022 [37] evaluation set, respectively. Furthermore, the MIMO-TSVAD framework demonstrates strong robustness against lip-missing problems. In simulated scenarios with varying lip-missing degrees, it guarantees that the DERs of the audio-visual system are always no worse than the audio-only system. In the future, we will further improve the current approach regarding advanced lip extraction and clustering methods, better use of multi-channel audio, etc.
Acknowledgments
This research is funded in part by the National Natural Science Foundation of China (62171207), Yangtze River Delta Science and Technology Innovation Community Joint Research Project (2024CSJGG01100), Science and Technology Program of Suzhou City (SYC2022051) and Guangdong Science and Technology Plan (2023A1111120012). Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.